# RGE effects on new physics searches via gravitational waves

### v2

@@ 1 Introduction
In this work, we address this open question and demonstrate that future GW observations can remain highly sensitive to various dimension-six SMEFT operators, provided that other experiments, including future collider measurements, precisely determine the effects of the $(H^{\dagger}H)^{3}$ operator. We adopt the SMEFT as a benchmark framework and comprehensively investigate the dimension-six operators that can affect the Higgs effective potential through one-loop RGE corrections involving the Higgs self-couplings, the top Yukawa coupling, and the gauge couplings. We focus on a scenario in which the $(H^{\dagger}H)^{3}$ operator drives the SFO-EWPT, while other dimension-six SMEFT operators introduce small shifts in the Higgs potential. We evaluate the impact of these SMEFT effects on the GW spectrum and perform a Fisher matrix analysis at two different renormalization scales, $\overline{\mu}_{\rm PT}=2\pi T$ and $T/2$, to clarify the expected sensitivity of future GW observations and their dependence on the renormalization scale. Furthermore, we repeat the analysis while varying the bubble wall velocity $v_{b}$ to examine the associated theoretical uncertainties.
@@ 2.1 Preliminary
$\mathcal{O}_{HD}$ $\ \left(H^{\dagger}D_{\mu}H\right)^{*}\left(H^{\dagger}D_{\mu}H\right)$
@@
Table 1: The dimension-six operators in the SMEFT [12] that are potentially relevant to the SFO-EWPT are listed. These operators affect the Higgs potential through both tree-level contributions and one-loop RGE effects involving the Higgs self-couplings $m^{2}$ and $\lambda$, the top Yukawa coupling $Y_{t}$, and the gauge couplings $g_{s}$, $g$, and $g^{\prime}$.
@@ 2.2 Matching at one-loop level
where $m_{\phi},m_{W}$, and $m_{t}$ are the $\overline{{\rm MS}}$-masses of Higgs, $W$ boson, and top quark, respectively, $g$ is the $\overline{{\rm MS}}$-$SU(2)_{L}$ gauge coupling, $M_{h},M_{W}$, and $M_{t}$ are pole masses of Higgs, $W$ boson, and top quark, respectively. When the tree-level gauge coupling $g_{0}$ is determined from the measurement of $G_{F}$, $g_{0}$ is expressed as $g_{0}^{2}:=4\sqrt{2}G_{F}M^{2}_{W}-8M^{2}_{W}C_{HW0}+2M_{W}^{2}\left[\left((C_{ll0})_{\mu ee\mu}+(C_{ll0})_{e\mu\mu e}\right)-2\left((C^{(3)}_{Hl0})_{ee}+(C^{(3)}_{Hl0})_{\mu\mu}\right)\right]$, where indices of 0 are used to explicitly indicate that the corresponding contributions arise at tree level. Also, $\Pi_{h,W}$, and $\Sigma_{t}$ are one-loop self-energies, and $\delta g^{2}$ is one-loop correction to the $SU(2)_{L}$ gauge coupling. The SMEFT operators listed in Table 1 potentially contribute to the one-loop correction appearing on the right-hand side of eqs. (2.20)(2.23) as follows:
@@
where the indices SM and SMEFT indicate the contributions from the SM and the SMEFT, respectively. The SM contributions [70] are listed in Appendix A. From eqs. (2.20)(2.23), the $\overline{{\rm MS}}$-parameters, e.g., eqs. (2.14) and (2.15), are improved at the one-loop level and are expressed in terms of the physical parameters. See Appendix B for the details. At the $\overline{\rm MS}$ scale $\overline{\mu}=M_{Z}$, the $\overline{\rm MS}$ parameters are matched at the one-loop level using eqs (B.1)(B.5). These parameters are evolved from $\overline{\mu}=M_{Z}$ to a typical scale of the SFO-EWPT using the RGEs, taking into account the effects of dimension-six SMEFT operators. The consequences of these SMEFT effects on the SFO-EWPT will be discussed in Section 3.
Now, we provide expressions of the one-loop SMEFT effects on eqs. (2.24)(2.27), and eqs. (B.4)(B.5). In what follows, we will examine three categories of one-loop effects within the SMEFT. These categories include corrections induced by Higgs self-couplings, associated with $m^{2}$ and $\lambda$; corrections arising from gauge couplings, including $g_{s}$, $g$, and $g^{\prime}$; and corrections induced by the top Yukawa coupling $Y_{t}$. Focusing on the logarithmic terms, we list the corresponding corrections in Tables 310 for the three types of self-energies: $\Pi_{h}$, $\Pi_{W}$, and $\Sigma_{t}$, as well as for the $SU(2)_{L}$ gauge coupling $g$, and the four SMEFT Wilson coefficients $c_{H,{\rm kin}}$, $C_{H}$, $C_{HW}$, and $C_{uH}$. While the logarithmic terms are listed in these tables for reference, the SMEFT RGE running is evaluated numerically in the analysis of the SFO-EWPT. See subsection 2.5 for the details of the procedures to evaluate the SFO-EWPT. In these Tables, the SMEFT operator one-loop correction normalized by $\frac{(4\pi)^{2}}{\ln\left(\overline{\mu}/M_{W}\right)}\times C_{\rm SMEFT}^{-1}$ is shown for each SMEFT Wilson coefficient $C_{\rm SMEFT}$. We showed these corrections by introducing parameters $h:=M_{h}/M_{W},t:=M_{t}/M_{W},z:=M_{Z}/M_{W}$, and $s:=g_{s}/\left(2M_{W}/v_{0}\right)$. As studied in Ref. [48], although these one-loop corrections have a small effect on the Higgs potential, they can still be measurable in GW observations once the SFO-EWPT occurs, e.g., via the SMEFT $(H^{\dagger}H)^{3}$ operator. Prescriptions to measure these effects will be explained in Section 3.
@@
$\mathcal{O}_{H}$ $-36M^{2}_{W}v_{0}^{2}\frac{1}{h^{2}}\left(-2+h^{4}+4t^{4}-z^{4}\right)$
@@
$\mathcal{O}_{HD}$ $-M^{4}_{W}\left(13h^{4}+3h^{2}\left(-2+2t^{2}+z^{2}\right)+12\left(-1+2t^{4}-2z^{4}\right)\right)$
$\mathcal{O}_{HWB}$ $-24M^{4}_{W}\sqrt{z^{2}-1}\left(h^{2}-3z^{2}\right)$
$\mathcal{O}_{HB}$ $-24M^{4}_{W}\left(h^{2}-3z^{2}\right)\left(z^{2}-1\right)$
@@
$(\mathcal{O}_{uH})_{33}$ $-12\sqrt{2}M^{3}_{W}v_{0}t\left(h^{2}-6t^{2}\right)$
@@
$(\mathcal{O}_{Hl}^{(3)})_{11}$ $-6M^{4}_{W}h^{2}\left(-2+h^{2}+2t^{2}-z^{2}\right)$
$(\mathcal{O}_{Hl}^{(3)})_{22}$ $-6M^{4}_{W}h^{2}\left(-2+h^{2}+2t^{2}-z^{2}\right)$
@@
$\mathcal{O}_{H\Box}$ $-\frac{4}{3}M^{4}_{W}\frac{1}{h^{2}}\left(-10h^{2}+15h^{4}+18\left(2-4t^{4}+z^{4}\right)\right)$
@@
$\mathcal{O}_{HB}$ $-24M^{4}_{W}\frac{z^{2}}{h^{2}}\left(-1+z^{2}\right)$
$\mathcal{O}_{HW}$ $-8M^{4}_{W}\left(-2+\frac{12}{h^{2}}+h^{2}-\frac{12t^{4}}{h^{2}}+\frac{3z^{2}}{h^{2}}+\frac{3z^{4}}{h^{2}}\right)$
$(\mathcal{O}_{uH})_{33}$ $-24\sqrt{2}M^{3}_{W}v_{0}\frac{t^{3}}{h^{2}}$
@@
$\mathcal{O}_{W}$ $-\frac{240M^{5}_{W}}{v_{0}}$
$(\mathcal{O}_{uW})_{33}$ $-24\sqrt{2}M^{4}_{W}t$
$(\mathcal{O}_{ll})_{2112}$ $-\frac{M_{W}^{4}}{3}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(20+18t^{2}-9z^{2})+18(2-4t^{4}+z^{4})\right)$
$(\mathcal{O}_{ll})_{1221}$ $-\frac{M^{4}_{W}}{3}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(20+18t^{2}-9z^{2})+18(2-4t^{4}+z^{4})\right)$
@@
$\mathcal{O}_{H\Box}$ $-2M^{2}_{W}\frac{1}{h^{2}}\left(5h^{4}-3h^{2}t^{2}+6\left(2-4t^{4}+z^{4}\right)\right)$
@@
$\mathcal{O}_{HB}$ $-12M^{2}_{W}\frac{1}{h^{2}}z^{2}\left(-1+z^{2}\right)$
$\mathcal{O}_{HW}$ $-12M^{2}_{W}\frac{1}{h^{2}}\left(2+z^{2}\right)$
@@
$(\mathcal{O}_{Hu})_{33}$ $-2M^{2}_{W}\left(-4+t^{2}+z^{2}\right)$
@@
$(\mathcal{O}_{qu}^{(1)})_{3333}$ $-8M^{2}_{W}t^{2}$
$(\mathcal{O}_{qu}^{(8)})_{3333}$ $-\frac{32}{3}M^{2}_{W}t^{2}$
$\mathcal{O}_{HG}$ $-64M_{W}^{2}s^{2}$
$\mathcal{O}_{H\widetilde{G}}$ $-64iM_{W}^{2}s^{2}$
$\mathcal{O}_{H\widetilde{W}}$ $-18iM^{2}_{W}$
$\mathcal{O}_{H\widetilde{B}}$ $-\frac{34}{3}iM^{2}_{W}\left(-1+z^{2}\right)$
@@
$\left(\mathcal{O}_{ll}\right)_{1221}$ $-\frac{M_{W}^{2}}{6}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(-8-9t^{2}+8z^{2})+18\left(2-4t^{4}+z^{4}\right)\right)$
$\left(\mathcal{O}_{ll}\right)_{2112}$ $-\frac{M_{W}^{2}}{6}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(-8-9t^{2}+8z^{2})+18\left(2-4t^{4}+z^{4}\right)\right)$
@@
$\left(\mathcal{O}_{ll}\right)_{1221}$ $-\frac{38}{3}M_{W}^{2}$
$\left(\mathcal{O}_{ll}\right)_{2112}$ $-\frac{38}{3}M_{W}^{2}$
@@
$\mathcal{O}_{W}$ $-\frac{120M_{W}^{3}}{v_{0}^{3}}$
$\left(\mathcal{O}_{uW}\right)_{33}$ $-\frac{12\sqrt{2}M_{W}^{2}}{v_{0}^{2}}t$
@@
$\mathcal{O}_{H\Box}$ $-\frac{2\sqrt{2}M_{W}^{3}}{3v^{3}_{0}}t\left(-20+3h^{2}+18t^{2}\right)$
@@
$\mathcal{O}_{HWB}$ $-\frac{4\sqrt{2}M_{W}^{3}}{v^{3}_{0}}t\sqrt{-1+z^{2}}$
@@
$\left(\mathcal{O}_{uG}\right)_{33}$ $-\frac{128M_{W}^{3}}{v^{3}_{0}}st^{2}$
$\left(\mathcal{O}_{uB}\right)_{33}$ $-\frac{8M_{W}^{3}}{v^{3}_{0}}\sqrt{-1+z^{2}}\left(-8+5t^{2}+5z^{2}\right)$
@@
$\mathcal{O}_{H\widetilde{W}B}$ $-i\frac{4\sqrt{2}M_{W}^{3}}{v^{3}_{0}}t\sqrt{-1+z^{2}}$
@@
$\left(\mathcal{O}_{Hq}^{(1)}\right)_{33}$ $-\frac{2\sqrt{2}M_{W}^{3}}{v^{3}_{0}}t\left(8+h^{2}+2t^{2}-8z^{2}\right)$

### v4

@@ 1 Introduction
In this work, we address this open question and demonstrate that future GW observations can remain highly sensitive to various dimension-six SMEFT operators, provided that other experiments, including future collider [70, 71], precisely determine the effects of the $(H^{\dagger}H)^{3}$ operator. We adopt the SMEFT as a benchmark framework and comprehensively investigate the dimension-six operators that can affect the Higgs effective potential through one-loop RGE corrections involving the Higgs self-couplings, the top Yukawa coupling, and the gauge couplings. We focus on a scenario in which the $(H^{\dagger}H)^{3}$ operator drives the SFO-EWPT, while other dimension-six SMEFT operators introduce small shifts in the Higgs potential. We evaluate the impact of these SMEFT effects on the GW spectrum and perform a Fisher matrix analysis at two different renormalization scales, $\overline{\mu}_{\rm PT}=2\pi T$ and $T/2$, to clarify the expected sensitivity of future GW observations and their dependence on the renormalization scale. Furthermore, we repeat the analysis while varying the bubble wall velocity $v_{b}$ to examine the associated theoretical uncertainties.
@@ 2.1 Preliminary
Table 1: The dimension-six operators in the SMEFT [12] that are potentially relevant to the SFO-EWPT are listed. These operators affect the Higgs potential through both tree-level contributions and one-loop RGE effects involving the Higgs self-couplings $m^{2}$ and $\lambda$, the top Yukawa coupling $Y_{t}$, and the gauge couplings $_{s}, gg$, and $g^{\prime}$.
@@ 2.2 Matching at one-loop level
where $m_{\phi},m_{W}$, and $m_{t}$ are the $\overline{{\rm MS}}$-masses of Higgs, $W$ boson, and top quark, respectively, $g$ is the $\overline{{\rm MS}}$-$SU(2)_{L}$ gauge coupling, $M_{h},M_{W}$, and $M_{t}$ are pole masses of Higgs, $W$ boson, and top quark, respectively. When the tree-level gauge coupling $g_{0}$ is determined from the measurement of $G_{F}$, $g_{0}$ is expressed as $g_{0}^{2}:=4\sqrt{2}G_{F}M^{2}_{W}-8M^{2}_{W}C_{HW0}+2M_{W}^{2}\left[\left((C_{ll0})_{\mu ee\mu}+(C_{ll0})_{e\mu\mu e}\right)-2\left((C^{(3)}_{Hl0})_{ee}+(C^{(3)}_{Hl0})_{\mu\mu}\right)\right]$, where indices of 0 are used to explicitly indicate that the corresponding contributions arise at tree level. Also, $\Pi_{h,W}$, and $\Sigma_{t}$ are one-loop self-energies, and $\delta g^{2}$ is one-loop correction to the $SU(2)_{L}$ gauge coupling. The SMEFT operators listed in Table 1 potentially contribute to the one-loop correction appearing on the right-hand side of eqs. (2.20)–(2.23) as follows:
@@
where the indices SM and SMEFT indicate the contributions from the SM and the SMEFT, respectively. The SM contributions [72] are listed in Appendix A. From eqs. (2.20)–(2.23), the $\overline{{\rm MS}}$-parameters, e.g., eqs. (2.14) and (2.15), are improved at the one-loop level and are expressed in terms of the physical parameters. See Appendix B for the details. At the $\overline{\rm MS}$ scale $\overline{\mu}=M_{Z}$, the $\overline{\rm MS}$ parameters are matched at the one-loop level using eqs (B.1)–(B.5). These parameters are evolved from $\overline{\mu}=M_{Z}$ to a typical scale of the SFO-EWPT using the RGEs, taking into account the effects of dimension-six SMEFT operators. The consequences of these SMEFT effects on the SFO-EWPT will be discussed in Section 3.
Now, we provide expressions of the one-loop SMEFT effects on eqs. (2.24)–(2.27), and eqs. (B.4)–(B.5). In what follows, we will examine three categories of one-loop effects within the SMEFT. These categories include corrections induced by Higgs self-couplings, associated with $m^{2}$ and $\lambda$; corrections arising from gauge couplings, including $_{s}, gg$, and $g^{\prime}$; and corrections induced by the top Yukawa coupling $Y_{t}$. Focusing on the logarithmic terms, we list the corresponding corrections in Tables 310 for the three types of self-energies: $\Pi_{h}$, $\Pi_{W}$, and $\Sigma_{t}$, as well as for the $SU(2)_{L}$ gauge coupling $g$, and the four SMEFT Wilson coefficients $c_{H,{\rm kin}}$, $C_{H}$, $C_{HW}$, and $C_{uH}$. While the logarithmic terms are listed in these tables for reference, the SMEFT RGE running is evaluated numerically in the analysis of the SFO-EWPT. See subsection 2.5 for the details of the procedures to evaluate the SFO-EWPT. In these Tables, the SMEFT operator one-loop correction normalized by $\frac{(4\pi)^{2}}{\ln\left(\overline{\mu}/M_{W}\right)}\times C_{\rm SMEFT}^{-1}$ is shown for each SMEFT Wilson coefficient $C_{\rm SMEFT}$. We showed these corrections by introducing parameters $h:=M_{h}/M_{W},t:=M_{t}/M_{W},z:=M_{Z}/M_{W}$, and $s:=g_{s}/\left(2M_{W}/v_{0}\right)$. As studied in Ref. [48], although these one-loop corrections have a small effect on the Higgs potential, they can still be measurable in GW observations once the SFO-EWPT occurs, e.g., via the SMEFT $(H^{\dagger}H)^{3}$ operator. Prescriptions to measure these effects will be explained in Section 3.
@@
$\mathcal{O}_{H}$ $36M^{2}_{W}v_{0}^{2}\frac{1}{h^{2}}\left(-2+h^{4}+4t^{4}-z^{4}\right)$
@@
$\mathcal{O}_{HD}$ $M^{4}_{W}\left(13h^{4}+3h^{2}\left(-2+2t^{2}+z^{2}\right)+12\left(-1+2t^{4}-2z^{4}\right)\right)$
$\mathcal{O}_{HWB}$ $24M^{4}_{W}\sqrt{z^{2}-1}\left(h^{2}-3z^{2}\right)$
$\mathcal{O}_{HB}$ $24M^{4}_{W}\left(h^{2}-3z^{2}\right)\left(z^{2}-1\right)$
@@
$(\mathcal{O}_{uH})_{33}$ $12\sqrt{2}M^{3}_{W}v_{0}t\left(h^{2}-6t^{2}\right)$
@@
$(\mathcal{O}_{Hl}^{(3)})_{11}$ $6M^{4}_{W}h^{2}\left(-2+h^{2}+2t^{2}-z^{2}\right)$
$(\mathcal{O}_{Hl}^{(3)})_{22}$ $6M^{4}_{W}h^{2}\left(-2+h^{2}+2t^{2}-z^{2}\right)$
@@
$\mathcal{O}_{H\Box}$ $\frac{4}{3}M^{4}_{W}\frac{1}{h^{2}}\left(-10h^{2}+15h^{4}+18\left(2-4t^{4}+z^{4}\right)\right)$
@@
$\mathcal{O}_{HB}$ $24M^{4}_{W}\frac{z^{2}}{h^{2}}\left(-1+z^{2}\right)$
$\mathcal{O}_{HW}$ $8M^{4}_{W}\left(-2+\frac{12}{h^{2}}+h^{2}-\frac{12t^{4}}{h^{2}}+\frac{3z^{2}}{h^{2}}+\frac{3z^{4}}{h^{2}}\right)$
$(\mathcal{O}_{uH})_{33}$ $24\sqrt{2}M^{3}_{W}v_{0}\frac{t^{3}}{h^{2}}$
@@
$\mathcal{O}_{W}$ $\frac{240M^{5}_{W}}{v_{0}}$
$(\mathcal{O}_{uW})_{33}$ $24\sqrt{2}M^{4}_{W}t$
$(\mathcal{O}_{ll})_{2112}$ $\frac{M_{W}^{4}}{3}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(20+18t^{2}-9z^{2})+18(2-4t^{4}+z^{4})\right)$
$(\mathcal{O}_{ll})_{1221}$ $\frac{M^{4}_{W}}{3}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(20+18t^{2}-9z^{2})+18(2-4t^{4}+z^{4})\right)$
@@
$\mathcal{O}_{H\Box}$ $2M^{2}_{W}\frac{1}{h^{2}}\left(5h^{4}-3h^{2}t^{2}+6\left(2-4t^{4}+z^{4}\right)\right)$
@@
$\mathcal{O}_{HB}$ $12M^{2}_{W}\frac{1}{h^{2}}z^{2}\left(-1+z^{2}\right)$
$\mathcal{O}_{HW}$ $12M^{2}_{W}\frac{1}{h^{2}}\left(2+z^{2}\right)$
@@
$(\mathcal{O}_{Hu})_{33}$ $2M^{2}_{W}\left(-4+t^{2}+z^{2}\right)$
@@
$(\mathcal{O}_{qu}^{(1)})_{3333}$ $8M^{2}_{W}t^{2}$
$(\mathcal{O}_{qu}^{(8)})_{3333}$ $\frac{32}{3}M^{2}_{W}t^{2}$
$\mathcal{O}_{HG}$ $64M_{W}^{2}s^{2}$
$\mathcal{O}_{H\widetilde{G}}$ $64iM_{W}^{2}s^{2}$
$\mathcal{O}_{H\widetilde{W}}$ $18iM^{2}_{W}$
$\mathcal{O}_{H\widetilde{B}}$ $\frac{34}{3}iM^{2}_{W}\left(-1+z^{2}\right)$
@@
$\left(\mathcal{O}_{ll}\right)_{1221}$ $\frac{M_{W}^{2}}{6}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(-8-9t^{2}+8z^{2})+18\left(2-4t^{4}+z^{4}\right)\right)$
$\left(\mathcal{O}_{ll}\right)_{2112}$ $\frac{M_{W}^{2}}{6}\frac{1}{h^{2}}\left(9h^{4}+h^{2}(-8-9t^{2}+8z^{2})+18\left(2-4t^{4}+z^{4}\right)\right)$
@@
$\left(\mathcal{O}_{ll}\right)_{1221}$ $\frac{38}{3}M_{W}^{2}$
$\left(\mathcal{O}_{ll}\right)_{2112}$ $\frac{38}{3}M_{W}^{2}$
@@
$\mathcal{O}_{W}$ $\frac{120M_{W}^{3}}{v_{0}^{3}}$
$\left(\mathcal{O}_{uW}\right)_{33}$ $\frac{12\sqrt{2}M_{W}^{2}}{v_{0}^{2}}t$
@@
$\mathcal{O}_{H\Box}$ $\frac{2\sqrt{2}M_{W}^{3}}{3v^{3}_{0}}t\left(-20+3h^{2}+18t^{2}\right)$
@@
$\mathcal{O}_{HWB}$ $\frac{4\sqrt{2}M_{W}^{3}}{v^{3}_{0}}t\sqrt{-1+z^{2}}$
@@
$\left(\mathcal{O}_{uG}\right)_{33}$ $\frac{128M_{W}^{3}}{v^{3}_{0}}st^{2}$
$\left(\mathcal{O}_{uB}\right)_{33}$ $\frac{8M_{W}^{3}}{v^{3}_{0}}\sqrt{-1+z^{2}}\left(-8+5t^{2}+5z^{2}\right)$
@@
$\mathcal{O}_{H\widetilde{W}B}$ $i\frac{4\sqrt{2}M_{W}^{3}}{v^{3}_{0}}t\sqrt{-1+z^{2}}$
@@
$\left(\mathcal{O}_{Hq}^{(1)}\right)_{33}$ $\frac{2\sqrt{2}M_{W}^{3}}{v^{3}_{0}}t\left(8+h^{2}+2t^{2}-8z^{2}\right)$